### v1

# Enhancing Group Fairness in Federated Learning through Personalization

Abstract
Personalized Federated Learning (FL) algorithms collaboratively train customized models for each client, enhancing the accuracy of the learned models on the client’s local data (e.g., by clustering similar clients, or by fine-tuning models locally). In this paper, we investigate the impact of such personalization techniques on the group fairness of the learned models, and show that personalization can also lead to improved (local) fairness as an unintended benefit. We begin by illustrating these benefits of personalization through numerical experiments comparing two classes of personalized FL algorithms (clustering and fine-tuning) against a baseline FedAvg algorithm, elaborating on the reasons behind improved fairness using personalized FL, and then providing analytical support. Motivated by these, we further propose a new, Fairness-aware Federated Clustering Algorithm, Fair-FCA, in which clients can be clustered to obtain a (tuneable) fairness-accuracy tradeoff. Through numerical experiments, we demonstrate the ability of Fair-FCA to strike a balance between accuracy and fairness at the client level.
1 Introduction
Federated Learning (FL) has emerged as a pivotal paradigm for collaboratively training a shared machine learning model across distributed datasets/clients in a privacy-preserving manner (Kairouz et al., 2021). Shared, global models learned through FL can (potentially) outperform standalone models (those trained individually by each client in the absence of collaboration), owing to their ability to effectively aggregate knowledge from diverse clients. Popular FL algorithms such as FedAvg (McMahan et al., 2017) have been found to perform well, particularly when clients have independent and identically distributed (iid) datasets. However, real-world applications often involve heterogeneous datasets across clients, a situation where the performance of FedAvg can considerably deteriorate (Li et al., 2019b). Moreover, the standard FL approach yields a single global model, lacking customization for each client. Consequently, clients with heterogeneous datasets may experience low local accuracy (Li et al., 2020; Tan et al., 2022; Karimireddy et al., 2020). To address these limitations, a spectrum of personalized FL techniques has been proposed, to enhance the local accuracy of the learned models, while keeping some of the benefits of collaborative learning.
In addition to handling issues of data heterogeneity, training FL models that can uphold societal values, as formalized through notions of algorithmic (group) fairness (Barocas et al., 2019), is an important consideration. Consider as an example the scenario where FL is used to train a foundation model/LLM on local datasets, contributed by diverse participants from different regions and communities (Kuang et al., 2023). Without careful oversight, the resulting model will favor language and content preferred by the majority contributors, often disregarding the unique linguistic nuances and cultural contexts of minority groups (Durmus et al., 2023). As another example, existing research (Kirdemir et al., 2021) finds structural and systemic bias in YouTube video recommendation systems. In this case, the collaboratively trained recommendation model begins to favor certain perspectives over others, inadvertently reinforcing biases present in the data. As a result, users are fed information that perpetuates stereotypes, causing harm to underrepresented communities. These examples motivate the development of FL algorithms that can meet certain fairness criteria.
In this paper, we establish an alignment between these two considerations in the developing FL algorithms: personalization can also improve fairness. That is, we show that personalization techniques (which are designed to improve the local accuracy of the learned models) can also enhance algorithmic fairness (i.e., they can reduce the disparities in how the resulting model treats individuals from different protected groups, once the model is deployed by the client).
We begin by illustrating this alignment between personalization and fairness through numerical experiments comparing the average local fairness achieved by two personalized FL algorithms, Clustered-FL (Ghosh et al., 2020) and MAML-FL (Fallah et al., 2020), against FedAvg and standalone learning (where each client learns by itself, and there is no collaboration), under three notions of fairness (Statistical Parity, Equality of Opportunity and Equalized Odds), and using both synthetic and real-world data (different prediction tasks in the “Retiring Adult” dataset (Ding et al., 2021)). Based on these experiments, we highlight two potential factors driving the alignment:
      1. Statistical advantages of diverse data. Collaborative learning algorithms in essence have access to more “information” (data) from diverse clients. They can therefore enhance both accuracy and fairness, especially when dealing with imbalanced datasets, as in these scenarios issues of algorithmic unfairness can be attributed to the under-representation of samples from the disadvantaged group in the data. Further, personalized algorithms can outperform non-personalized collaborative learning (e.g. FedAvg), as (we posit) they are less prone to overfitting to the majority group’s data.
      2. Computational advantages due to alignments in local accuracy and fairness. Moreover, we identify cases where the sample distribution within clients is such that improving accuracy also promotes fairness. More intuitively, consider the concept of clustered FL, a personalized FL technique, where clients are organized into groups based on similarities in model performance (a proxy for data similarities). We argue that this clustering can be seen as a means to foster a model that is both more precise and more equitable, as it effectively treats information originating from other clusters as noise, which, if left unaddressed, would have led to model divergence, potentially compromising both cluster-specific performance and fairness. Building on this intuition, we also provide analytical support of the alignment between personalization and fairness (Propositions 1 and 2).
To the best of our knowledge, this work is the first to identify the (unintended) fairness benefits of personalization techniques in FL. Most existing works primarily focus on refining algorithms to attain either improved local accuracy (the personalized FL literature), or, enhanced (global) fairness (the fair FL literature); we discuss related work from each direction in Section 2. In contrast, our work examines the influence of existing personalization techniques on (local) fairness, and points to inherent features of collaborative learning and personalization that can advance fairness.
Additionally, prior work points to the challenge of balancing the trade-off between fairness and accuracy of an algorithm, as enhancing fairness often comes at the cost of accuracy degradation (Gu et al., 2022; Ezzeldin et al., 2023; Chien & Danks, 2023). In contrast, we identify instances that FL in general, and personalized FL in particular, can improve both accuracy and fairness compared to standalone learning. Furthermore, inspired by our findings, we propose a new, Fairness-aware Federated Clustering Algorithm, Fair-FCA. Our algorithm is an extension of the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020), by modifying the cluster identity assignment step to encompass a fairness criterion. Given our finding that personalization can offer dual benefits, we find that this modification can be used to strike a better balance between accuracy and fairness at the client level.
Summary of findings and contributions.
1.
We conduct extensive numerical experiments, under different notions of fairness, and using both real-world and synthetic data, to show that personalization in FL (which is intended to improve local accuracy) can also improve local fairness as an unintended benefit (Section 4). Accordingly, we highlight the potential statistical and computational reasons leading to this alignment.
2. We analytically show that, under certain conditions, personalization via clustering can lead to improved local fairness compared to a (non-personalized) FedAvg model (Propositions 1 and 2).
3. We propose a Fairness-aware Federated Clustering Algorithm, Fair-FCA, an extension of the IFCA federated clustering algorithm (Ghosh et al., 2020), which takes fairness into account when (iteratively) determining clients’ cluster memberships (Algorithm 1), and show that it can provide desired fairness-accuracy trade-offs by adjusting a hyperparameter (Section 6).
2 Related works
Our work is at the intersection of two literatures: personalized FL, and achieving fairness in FL. Here, we discuss the most closely related works; we review additional related work in Appendix B.
Personalized FL. Existing literature can be categorized based on how personalization is achieved. In this paper, we investigate personalization through clustering and local fine-tuning.
Clustering: Mansour et al. (2020) use a hypothesis-based clustering approach by minimizing the sum of loss over all clusters. Sattler et al. (2020) use the idea that cosine similarity between weight updates of different clients is highly indicative of the similarity of data distribution. Nardi et al. (2022) use a decentralized learning idea by exchanging local models with other clients to find the neighbor/group which has a high accuracy even using other clients’ models. Zheng et al. (2022) learn a weighted and directed graph that indicates the relevance between clients. Ghosh et al. (2020) use a distributed learning idea by broadcasting all clients models to others, and collecting back the cluster identity from clients who can identify good performance when using others’ models.
Local fine-tuning: Fallah et al. (2020) propose using a Model Agnostic Meta Learning (MAML) framework, where clients run additional local gradient steps to personalize the global model. Arivazhagan et al. (2019); Jiang & Lin (2022) propose using deep learning models with a combination of feature extraction layers (base) and global/local head (personalization). Jiang & Lin (2022), inspired by Arivazhagan et al. (2019), further consider robustifying against distribution shifts.
In this paper, we will consider the Ghosh et al. (2020) Clustered-FL (IFCA) algorithm from the clustering category, and the Fallah et al. (2020) MAML-FL algorithm from the local fine-tuning category. Further, our proposed algorithm, Fair-FCA, belongs to the clustering category.
Fairness in FL. This literature, surveyed recently in Shi et al. (2023); Rafi et al. (2023), can also be categorized depending on the adopted notion of fairness. Notably, several works (e.g., (Li et al., 2019a, 2021; Zhang et al., 2021; Wang et al., 2021)) on fair FL consider a notion of performance fairness, assessing whether the learned model(s) achieve uniform accuracy across all clients. In contrast, we focus on notions of social (group) fairness (also studied in, e.g., (Abay et al., 2020; Zhang et al., 2020; Du et al., 2021; Gálvez et al., 2021; Zeng et al., 2021; Ezzeldin et al., 2023)), which compare statistics such as selection rate, true positive rate, or false positive rate, across different demographic/protected groups, to identify any bias in how humans will be treated once the learned models are deployed. Specifically, Abay et al. (2020) propose pre-processing (reweighting with differential privacy) and in-processing (adding fairness-aware regularizer) methods to mitigate biases while protecting the data privacy. Zhang et al. (2020) propose a reinforcement-learning based, fair FL framework, in which the reward is aligned with maximizing the global accuracy while minimizing the discrimination index between groups. Du et al. (2021) propose a fairness-aware FL framework to train a globally fair model through a kernel reweighting technique on both loss function and fairness constraints. Gálvez et al. (2021) introduce an algorithm to enforce group fairness in FL by extending the modified method of differential multiplier to empirical risk minimization with fairness constraints. Ezzeldin et al. (2023) propose fairness-aware aggregation in FL by adjusting the model aggregation weights to prioritize clients whose local fairness is lower than the global fairness. Similarly, Zeng et al. (2021) update model aggregation weights, but by iteratively solving a bi-level optimization problem similar to the centralized fair learning algorithm.
In contrast to all these works, we do not aim to impose a fairness constraint, but show that improved (local) fairness (and a better fairness-accuracy tradeoff) can be achieved either by personalization alone, or by grouping clients into clusters based on a fairness-aware cluster assignment metric.
3 Problem setting and preliminaries
We study a federated learning setting consisting of $n$ clients. The clients can (potentially) be categorized into two clusters, denoted $c\in\{\mathcal{C}_{\alpha},\mathcal{C}_{\beta}\}$, based on similarities in their data distributions. (We consider $K\geq 2$ clusters in Section 6.) Let $p$ denote the fraction of clients belonging to cluster $\mathcal{C}_{\alpha}$.
Each client $i$ is tasked with a binary classification problem. The client’s dataset consists of samples $z=(x,y,g)$, where $x\in\mathbb{R}$ is a feature, $y\in\{0,1\}$ is the true label of the sample, and $g\in\{a,b\}$ is a demographic group membership based on a protected attribute (e.g., race, gender, or age). Client $i$’s samples are drawn independently at random from joint feature-label-group distributions with pdfs $\{f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\}\}$; note that these distributions are cluster-dependent.
Standalone learning. Clients can use their local data to (independently) learn a threshold-based, binary classifiers $h_{\theta}(x):\mathbb{R}\rightarrow\{0,1\}$, [^1] Our analysis will assume one-dimensional features and threshold classifiers. The former can be viewed as the one-dimensional representation of multi-dimensional features obtained from the last layer outputs of a neural network. For the latter, existing works (Corbett-Davies et al., 2017; Raab & Liu, 2021) show that threshold classifiers are optimal when multi-dimensional features can be properly mapped into a one-dimensional space. under which samples with features $x\geq\theta$ are classified as label 1 (i.e., $\hat{y}(\theta)=1$). Clients choose these thresholds to minimize classification errors. Formally, consider a client $i$ from cluster $c$; let $r^{c}_{g}$ be the fraction of its samples that are from group $g$, and $\alpha^{y,c}_{g}$ be the fraction of its samples that are from group $g$ and have true label $y$. The client chooses its decision threshold ${\theta_{i}^{*}}$ to (empirically) solve the following optimization problem:
$\theta_{i}^{*}=\arg\min_{{\theta}}~{}~{}\sum_{g\in\{a,b\}}r^{c}_{g}\Big{(}\alpha^{1,c}_{g}\int_{-\infty}^{\theta}f^{1,c}_{g}(x)\mathrm{d}x+\alpha^{0,c}_{g}\int_{\theta}^{+\infty}f^{0,c}_{g}(x)\mathrm{d}x\Big{)}~{}.$ (1)
Federated learning. While clients can learn standalone models, Federated learning paradigms enable them to leverage the power of collaborative learning. This is exemplified by the FedAvg algorithm (McMahan et al., 2017), in which the objective is to learn a single global model $\theta$ that minimizes the average loss across all clients. Learning happens over rounds. During each communication round $t$, each client $i$ receives the latest global model $\theta^{t-1}$ from the server, performs a local update on this model based on its local data, and sends back its updated model $\theta^{t}_{i}$ to the server. The server then aggregates all the received local models (through a weighted sum) to create the new global model for the next round. While FedAvg can successfully learn a global model, if the clients are heterogeneous, the local loss of the shared global model may be high locally for each client. This has motivated the development of personalized FL algorithms, two types of which we consider in this work: clustering and local fine-tuning. The Clustered-FL algorithm of Ghosh et al. (2020) alternates between two steps: clustering/grouping clients based on model performance, and optimizing model parameters via gradient descent in each cluster separately, leading to distinct, cluster-specific models. Similarly, the MAML-FL algorithm of Fallah et al. (2020) takes an extra step beyond the FedAvg approach. In it, clients fine-tune the received global model by running additional gradient steps on it locally, to specialize it to their local data and reduce their local loss.
Assessing Fairness. Given the sensitive demographic information $g$ (e.g., race, gender, age) contained in each data point in our problem setting, we further assess the group fairness of the learned models. Different notions of group fairness have been proposed, each assessing a specific statistical disparity in how the model treats individuals from different groups. Here, we consider three notions of fairness: (1) Statistical Parity (SP) (Dwork et al., 2012; Ezzeldin et al., 2023), which compares the difference in selection rate between the two groups (i.e., $\Delta_{\texttt{SP}}(\theta):=|{Pr}(\hat{y}(\theta)=1|g=a)-{Pr}(\hat{y}(\theta)=1|g=b)|$); (2) Equality of Opportunity (EqOp) (Hardt et al., 2016), which compares the differences in true positive rates across groups (i.e., $\Delta_{\texttt{EqOp}}(\theta):=|Pr(\hat{y}(\theta)=1|g=a,y=1)-{Pr}(\hat{y}(\theta)=1|g=b,y=1)|)$; and (3) Equalized Odd (EO) (Hardt et al., 2016), which is set to either the true positive rate or false positive rate difference between the two protected groups, whichever larger (i.e., $\Delta_{\texttt{EO}}(\theta):=\max_{i\in\{0,1\}}|{Pr}(\hat{y}(\theta)=1|g=a,y=i)-{Pr}(\hat{y}(\theta)=1|g=b,y=i)|$).
Comparing different algorithms. To compare the different algorithms’ performance, we will contrast the (average of the) local accuracy of each client’s local model (which can be its standalone model, the global FedAvg model, a cluster-specific model, or a fine-tuned model) when applied to the client’s local data. We will similarly contrast the different algorithms in terms of the (average of) local $f$-fairness, evaluating each client’s local model on its local data under $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$.
4 Personalization can also improve fairness: numerical illustration
We begin by numerically illustrating the (unintended) fairness benefits of personalization, and elaborate on the reasons behind it. (We substantiate these with analytical support in Section 5.) We will compare the average local statistical parity (SP) fairness achieved by two personalized FL algorithms, Clustered-FL (Ghosh et al., 2020) and MAML-FL (Fallah et al., 2020), against FedAvg and standalone learning (where each client learns by itself, and there is no collaboration). Similar experiments supporting our findings under other notions of fairness (EqOp, EO) are given in Appendix E.
We present numerical experiments on the “Retiring Adult” dataset (Ding et al., 2021). The dataset consists of census data collected from 50 U.S. states, and Puerto Rico. In our context, each state is a client within the FL framework. Each data sample includes multi-dimensional features $x$ (e.g., age, education, citizenship), a true label $y$, and a protected attribute $g$ (race, sex). To provide a more clear comparison of the effects of personalization, we have manually scaled the feature set ($x$) by 60% for the states with IDs {1, 10, 20, 30, 40, 50}; this exacerbates the data heterogeneity. We focus on two binary classification tasks: Employment (ACSEmployment) and Income (ACSIncome), and employ a two-layer neural network for both tasks. Each client is provided with 1000 training samples and 2000 testing samples. The values reported are the averages from 5 runs. Further details regarding the dataset and models, and additional figures, can be found in Appendix C.
4.1 Imbalanced groups: statistical advantages of collaboration
We first consider the ACSEmployment task with “race” as the protected attribute. Fig 1(a) shows the fraction of samples in each group-label, from several states, highlighting an imbalance between samples from the White and Non-White groups. This is further evident in Figure 1(b), which shows that most states have only $\sim 10\%$ qualified (label 1) samples from the Non-White group, in contrast to $\sim 35\%$ qualified samples from the White group.
(a) Fraction of samples
(b) Normalized sample frequency
(c) Local accuracy vs. fairness gap
Figure 1: Experiments on the ACSEmployment task with imbalanced groups (race).
Fig 1(c) shows that all collaborative training algorithms (FedAvg, MAML-FL, Clustered-FL) achieve better local fairness (smaller gap) compared to standalone learning. This is due to the statistical benefits of collaboration: each client has limited samples in the non-White group, leading to poorly trained models with high local fairness gap (and low accuracy). In contrast, collaborative training in essence has access to more data, improving both metrics. For the same reason, the Clustered-FL algorithm, which partitions clients into two clusters, has (slightly) worse local fairness compared to FedAvg. In comparison, the MAML-FL algorithm, which effectively sees the global dataset (when training the global model that is later fine-tuned by each client), has better local fairness compared to FedAvg, indicating that personalization can improve both local accuracy (as intended) and local fairness (as a side benefit).
4.2 Better-balanced groups: computational advantages of collaboration
We next consider better-balanced data, to show advantages of collaborative and personalized training beyond the statistical benefits of (effectively) expanding training data. We again consider the ACSEmployment task, but now with “sex” as the protected attribute. Fig 2(a) shows that data samples are more evenly distributed across groups and labels in this problem. Figure 2(b) further confirms that clients exhibit similar sample fractions of label 1 individuals in male and female groups.
(a) Fraction of samples
(b) Normalized sample frequency
(c) Local accuracy vs. fairness gap
Figure 2: Experiments on the ACSEmployment task with better-balanced groups (sex).
We first notice from Fig 2(c) that all collaborative training algorithms still have better local fairness compared to standalone learning. Furthermore, we observe that Clustered-FL and MAML-FL improve both local accuracy and local fairness compared to FedAvg. This is due to the computational advantages of (personalized) collaborative learning: for each client, due to similarity of the data for the male and female groups (as seen in Figure 2(b)) the objective of maximizing local accuracy is aligned with reducing the local fairness gap. Therefore, collaboration improves local fairness, with personalization further enhancing the model’s local accuracy and therefore its fairness.
We also note that (local) accuracy and fairness may not necessarily be aligned. Our next experiment shows that personalization can still improve fairness in such tasks compared to non-personalized FedAvg, which (we posit) is driven by a combination of statistical and computational benefits. Specifically, we conduct experiments on another task, ACSIncome, with “sex” as the protected attribute. Fig 3(a) shows that for this task, the fraction of samples is comparable across groups for label 0 data, but differs for label 1 data. From Fig 3(c), we observe that this time, all collaborative training algorithms improve accuracy but have worse local fairness compared to standalone learning; this is because improving (local) accuracy is not aligned with fairness in this task. That said, we observe that the personalized FL algorithms slightly improve local fairness compared to FedAvg. We interpret this as the statistical advantage of (effectively) observing more label 1 data (as FedAvg does, too), combined with a computational advantage of not overfitting a global model to the majority label 0 data (unlike what FedAvg may be doing).
(a) Fraction of samples
(b) Normalized sample frequency
(c) Local accuracy vs. fairness gap
Figure 3: Experiments on the ACSIncome task with sex as the protected attribute.
4.3 Comparison with a fair Federated Learning algorithm
Figure 4: Comparing FedAvg, FairFed, Clustered-FL, MAML-FL, and standalone learning.
A natural question that may arise is why not utilize an existing fair FL algorithm to improve fairness, as these might offer superior fairness compared to a personalized FL algorithm. Indeed, if one only focuses on improving (global) fairness, choosing a dedicated fair FL algorithm could be the best choice. However, here we point to the additional (cost-free) local fairness improvements achieved through the incorporation of personalization in FL. Our findings also suggest that this leads to a more favorable trade-off between fairness and accuracy, inspiring our new algorithm in Section 6.
To show this, we consider the FairFed algorithm of Ezzeldin et al. (2023), which adjusts the global model aggregation weights according to a fairness metric with the goal of improving (global) fairness. We also consider an extension of the algorithm, FairFed-AF, which adjusts the weights in aggregation according to both accuracy and fairness metrics. The experiments in Fig. 4 are on the ACSEmployment task with sex as the protected attribute. We observe that fair FL algorithms can achieve the best fairness across all algorithms. However, they have worse local accuracy compared to other collaborative training algorithms, as they focus (at least partially) on improving fairness and not accuracy.
5 Personalization can also improve fairness: theoretical support
To support and validate our findings from the numerical experiments in Section 4, in this section, we analytically show that personalized Federated clustering algorithms (which cluster/group similar clients to improve their models’ local accuracy) can also lead to better local fairness, when compared to a (non-personalized) shared global model.
Recall from the setup of Section 3 that we assume the $n$ clients can be potentially grouped into two clusters, $C_{\alpha}$ and $C_{\beta}$, based on their data similarities in their data distributions $f^{y,c}_{g}(x)$, with a fraction $p$ of clients in cluster $C_{\alpha}$. For personalized learning, we consider a cluster-based FL algorithm where each cluster can learn its own optimal cluster-specific model $\theta^{*}_{i},~{}i\in\{\alpha,\beta\}$ (obtained after solving equation 1), and contrast that with the average optimal model $\theta^{*}_{G}$ that would be obtained if all $n$ clients collaboratively learn a shared global model. We then contrast the average local fairness $\Delta^{\alpha}_{f}(\theta)$ obtained for clients in cluster $C_{\alpha}$ under a personalized model $\theta^{*}_{\alpha}$ vs. a shared model $\theta^{*}_{G}$, for two notions of fairness: $f\in\{\texttt{EqOp},\texttt{SP}\}$.
We start with the EqOp (Equality of Opportunity) fairness constraint, which aims to equalize true positive rates (TPR) between the protected groups $a$ and $b$. The following proposition shows that if $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$ (i.e., the data heterogeneity is such that cluster $C_{\alpha}$ has a lower optimal threshold than $C_{\beta}$), then clients in cluster $C_{\alpha}$ can obtain better local fairness (in addition to better local accuracy) with their cluster-specific model compared to if they used a global model shared with clients in $C_{\beta}$.
Proposition 1 (Improved EqOp through clustering).
Assume $f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\},c\in\{C_{\alpha},C_{\beta}\}$, are unimodal distributions, with modes $m^{y,c}_{g}$ such that $m^{y,c}_{b}\leq m^{y,c}_{a},\forall i,c$, and $\alpha^{1,c}_{g}\geq\alpha^{0,c}_{g},\forall g,c$. If $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$, there exist a cluster size $\hat{p}$ such that for $p\geq\hat{p}$, we have $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})<\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$; that is, the global model is more unfair than the cluster-specific model for $C_{\alpha}$.
The proof is presented in Appendix D.1. Intuitively, clients in $C_{\alpha}$ are better off under their personalized model as, given the proposition’s conditions, an increase in the decision threshold (which happens when moving from $\theta^{*}_{i}$ to $\theta^{*}_{G}$) will decrease the TPR of the disadvantaged group $b$ (the one with a lower mode in its feature distribution) slower than that of the advantaged group $a$, increasing the fairness gap for clients in $C_{\alpha}$.
(a) SP
(b) EqOp
Figure 5: Fairness gap vs $\theta$.
We next consider SP (Statistical Parity) fairness, which assesses to disparity in the selection (positive classification) rate between the two protected groups. This is impacted by both the group $a$ vs. $b$ feature distributions as well as the label rates, rendering it more stringent than EqOp fairness. Figure 5 illustrates this by plotting the fairness gap vs. the decision threshold $\theta$ for SP vs. EqOp, showing that SP exhibits less structured changes as the decision threshold moves (e.g., due to the use of a global model). Therefore, to facilitate theoretical analysis, we further assume that $f^{y,c}_{g}(x)$ follows Gaussian distributions with equal variance $\sigma^{2}$ but different means $\mu^{y,c}_{g}$, such that $\mu^{0,c}_{b}\leq\mu^{0,c}_{a}\leq\mu^{1,c}_{b}\leq\mu^{1,c}_{a}$; the ordering is chosen so that label 0 samples have lower features than label 1 samples, and that for the same label, group $a$ samples have higher features than group $b$ samples (making group $a$ advantaged). We again find that then clients in cluster $C_{\alpha}$ can obtain better local SP-fairness (in addition to better local accuracy) with their cluster-specific model compared to if they joined in on a global model shared with clients in $C_{\beta}$.
Proposition 2 (Improved SP through clustering).
Assume $f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\},c\in\{C_{\alpha},C_{\beta}\}$, are Gaussian distributions with means $\mu^{0,c}_{b}\leq\mu^{0,c}_{a}\leq\mu^{1,c}_{b}\leq\mu^{1,c}_{a}$ and equal variance $\sigma^{2}$. Assume further that $\mu^{1,c}_{a}-\mu^{0,c}_{a}=\mu^{1,c}_{b}-\mu^{0,c}_{b}$, $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$, and that either $\alpha^{1,c}_{g}\geq\alpha^{0,c}_{g},\forall g$, or $r^{c}_{b}\geq r^{c}_{a},\forall c$ and $\alpha^{1,c}_{a}>\alpha^{0,c}_{b}>\alpha^{1,c}_{b}>\alpha^{0,c}_{a}$. Then, if $\alpha^{0,c}_{a}\exp(\frac{(\bar{\theta}-\mu^{0,c}_{a})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{0,c}_{a})-\alpha^{1,c}_{b}\exp(\frac{(\bar{\theta}-\mu^{1,c}_{b})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{1,c}_{b})\geq\alpha^{0,c}_{b}\exp(\frac{(\bar{\theta}-\mu^{0,c}_{b})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{0,c}_{b})-\alpha^{1,c}_{a}\exp(\frac{(\bar{\theta}-\mu^{1,c}_{a})^{2}}{-2\sigma^{2}})(\bar{\theta}-\mu^{1,c}_{a})$ holds, where $\bar{\theta}:=\tfrac{\mu^{1,c}_{a}+\mu^{0,c}_{b}}{2}$, there exist a $\hat{p}$ such that for $p\geq\hat{p}$, $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})<\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$.
A detailed proof is given in Appendix D.2. Proposition 2 assumes an equal distance between mean estimates (reflecting a uniform, systematic underestimation of group $b$ features); we relax this in Appendix E.7-E.10. Intuitively, the proposition states the following: when $\alpha^{1,c}_{g}>\alpha^{0,c}_{g}$, there are more label 1 data in both groups, and $\theta^{*}_{G}$ will pull $\theta^{*}_{\alpha}$ up to account for the label imbalance, resulting in a deterioration in both fairness and accuracy. Similarly, the other condition means that group $a$’s clients are majority label 1, while group $b$’s clients are majority label 0; then, $\theta^{*}_{\alpha}<\bar{\theta}$ if $r_{b}\geq r_{a}$, resulting in a higher fairness gap for clients in $\mathcal{C}_{\alpha}$ under $\theta^{*}_{G}$ for the same reason as the first condition.
6 Fair-FCA: A Fairness-aware Federated Clustering Algorithm
We have observed numerically (Section 4) and analytically (Section 5) that integrating personalization techniques (such as clustering) in the FL paradigm can provide a dual benefit by not only enhancing local accuracy as intended, but also improving (local) group fairness. Motivated by this, we propose a new Fairness-aware Federated Clustering Algorithm (Fair-FCA), which builds on existing clustered FL ideas to further integrate fairness considerations in cluster identity assignment.
6.1 The Fair-FCA algorithm
Our algorithm is an extension of the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020), by modifying the cluster identity assignment step to encompass a fairness criterion. Specifically, we consider a FL scenario involving a total of $n$ clients and $K$ disjoint clusters. Each client $i$ has $n_{i}$ samples $Z_{i}:=\{z_{ij}\}^{n_{i}}_{j=1}$. Let $f(z,\theta)$ be the loss function associated with data point $z$ under model $\theta$. Then, for client $i$, the empirical loss is given by $F(Z_{i},\theta):=\frac{1}{n_{i}}\sum_{j}f(z_{ij},\theta)$. Additionally, let $\Psi^{f}(Z_{i},\theta)$ represent the fairness under model $\theta$ for fairness metric $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$ assessed on a dataset $Z_{i}$ (i.e., $\Delta_{f}(\theta)$ empirically calculated on $Z_{i}$).
Our algorithm iterates over two steps: (1) cluster identity assignment, and (2) training of cluster-specific models. Specifically, let $\Theta^{t}_{k}$ denote cluster $k$’s model at time step $t$. The cluster identity for client $i$ at time $t$, denoted $c^{t}(i)$, is determined by:
$c^{t}(i)=\arg\min_{k\in[K]}\gamma F(Z_{i},\Theta^{t}_{k})+(1-\gamma)\Psi^{f}(Z_{i},\Theta^{t}_{k})$ (2)
Here, $\gamma$ is a hyperparameter that strikes a desired balance between accuracy and fairness. For $\gamma=1$, we recover the IFCA algorithm; for $\gamma=0$, we obtain a clustered FL algorithm that prioritizes (local) $f$-fairness only when grouping clients. For $10<\gamma<$ 1, we obtain clusters that provide each client with the best fairness-accuracy tradeoff among those attainable if the client were to join each cluster. Let $C^{t}_{k}$ be the set of clients whose cluster identity is $k$ at the end of this assignment process (i.e., $C^{t}_{k}=\{i\in[n]:c^{t}(i)=k\}$).
Once clients get assigned clusters, each client $i$ starts from its corresponding cluster model $\Theta^{t}_{c^{t}(i)}$, and locally runs gradient steps to update this model, such that $\theta_{i}^{t}=\Theta^{t}_{c^{t}(i)}-\eta\nabla_{\theta_{i}}F(Z_{i},\Theta^{t}_{c^{t}(i)})$. Then, the updated local models $\theta_{i}^{t}$ are sent to the central server, who uses these to update the cluster models to $\Theta^{t+1}_{1:K}$ by taking the weighted average of the local models of clients in corresponding clusters. Formally, $\Theta^{t+1}_{k}=\Theta^{t}_{k}-\sum_{i\in C^{t}_{k}}\frac{n_{i}}{\sum_{i}n_{i}}(\Theta^{t}_{k}-\theta^{t}_{i}),\forall k\in[K]$. The pseudo-code for Fair-FCA is shown in Algorithm 1.
Input: Number of clusters $K$, number of clients $n$, number of local updates $E$, cluster model initialization $\Theta_{1:K}$, learning rate $\eta$, fairness-accuracy tradeoff $\gamma$, fairness $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$.
Initialize: Start clusters $k\in[K]$ by randomly selecting one client for each.
while not converge do
      for client $i\in[n]$ do
             Find cluster identity:
                     $c(i)=\arg\min_{k\in[K]}\gamma F(Z_{i},\Theta_{k})+(1-\gamma)\Psi^{f}(Z_{i},\Theta_{k})$
             Initialize $\theta_{i}=\Theta_{c(i)}$
             Perform $E$ steps of local update
                     $\theta_{i}=\theta_{i}-\eta\nabla_{\theta_{i}}F(Z_{i},\theta_{i})$
             Upload $\theta_{i}$ to server
       end for
      Update the cluster model $\Theta_{1:K}$
               $\Theta_{k}=\Theta_{k}-\sum_{i\in C_{k}}\frac{n_{i}}{\sum_{i}n_{i}}(\Theta_{k}-\theta_{i})$
       Send new cluster models $\Theta_{1:K}$ to all clients
end while
Output: Cluster models $\Theta_{1:K}$, Cluster identity $c(i),\forall i\in[n]$
Algorithm 1 Fair-FCA: A Fairness-aware Federated Clustering Algorithm
6.2 Fairness-accuracy tradeoff using Fair-FCA
Figure 6: Acc. vs. SP-fairness
We begin by conducting a numerical experiment on a synthetic dataset, to illustrate the ability of Fair-FCA to strike a balance between fairness and accuracy. We consider a total of 8 clients that could (potentially) be clustered into two clusters. Among these, 6 clients (Client ID: 2,4,5,6,7,8) have similar data distributions, with 4 clients (Client ID: 4,6,7,8) sharing identical distributions across the two protected groups $a,b$ (low fairness gap). The remaining 2 clients (Client ID: 1,3) have different data distributions compared to the first six, but they also share identical distributions across the two protected groups. We consider Fair-FCA on this problem, with $f=\texttt{SP}$, and $\gamma\in\{0,0.3,0.5,0.8,1\}$.
Figure 6 shows that when $\gamma=1$, Fair-FCA prioritizes accuracy; by design, this is attained by grouping the 6 clients having similar data distributions together ({1,3} and {2,4,5,6,7,8}). Similarly, when $\gamma=0$, Fair-FCA focuses only on SP fairness, this time clustering clients that have identical distributions on the two protected groups together ({2,5} and {1,3,4,6,7,8}). Lastly, by setting $\gamma\in(0,1)$, we can effectively account for both accuracy and SP fairness when clustering: when $\gamma=0.3$, the clusters are {2,4,5} and {1,3,6,7,8}; when $\gamma=0.5$, the clusters are {2,4,5,6} and {1,3,7,8}; and when $\gamma=0.8$, the clusters are {2,4,5,6,7} and {1,3,8}.
6.3 Numerical experiments: comparing Fair-FCA vs. IFCA (Clustered-FL)
We now compare the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020) and our Fair-FCA algorithm with $\gamma=0.5$ and with $f=\texttt{SP}$, in terms of local accuracy and local group fairness. We proceed with the same experiment setting as in Section 4. (We also provide experiment results using the original Retiring Adult dataset without feature scaling in Appendix E.5, and experiments with EqOp as the fairness metric in Appendix E.3.)
Figure 7 verifies that under different datasets (ACSEmployment, ACSIncome) and different protected attributes (race, sex), Fair-FCA exhibits a fairness-accuracy tradeoff (improved fairness performance at the expense of accuracy degradation) consistent with the choice of $\gamma=0.5$ compared to the Clustered-FL algorithm.
(a) Employment (Race)
(b) Employment (Sex)
(c) Income (Sex)
Figure 7: Fair-FCA with different tasks and protected attributes ($f=\texttt{SP}$)
7 Conclusion
We have shown, both numerically and analytically, that there can be (unintended) fairness benefits to personalization in Federated Learning. We have identified potential statistical reasons (improved data diversity) and computational reasons (alignments between local accuracy and fairness while preventing overfitting) for this alignment. Motivated by these findings, we further proposed a Fairness-aware Federated Clustering Algorithm (Fair-FCA), which takes both the fairness and model performance into account when clustering clients. We find that this modification can lead to an improved fairness-accuracy tradeoff, (partly) as personalization technique offer dual benefits in terms of accuracy and fairness. Experiments under other types of personalized FL algorithms (beyond clustering and local fine-tuning), and extending our analytical findings (both for clustered FL algorithms, and to other classes of personalized FL methods) are main directions of future work.
Acknowledgments and Disclosure of Funding
This work is supported by Cisco Research, and by the National Science Foundation (NSF) program on Fairness in AI in collaboration with Amazon under Award No. IIS-2040800. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF, Amazon, or Cisco.
Appendix
Appendix A Limitations and extensions
Personalization techniques. Our numerical and analytical results primarily focus on the IFCA framework. Although we also evaluate the performance of the MAML FL algorithm in our experiments, additional experimental results using other types of local fine-tuning methods could be valuable to explore. Additionally, extending the analytical results for other clustered FL methods and other classes of personalized FL methods are main directions of future work.
FL setting. Our theoretical support for the idea that personalization can improve fairness assumes identical clients within two clusters. Extending this framework to include multiple clusters and account for client heterogeneity within each cluster is a straightforward extension. Additionally, allowing for client drop-out, a scenario not considered in the current FL setting, presents another avenue for exploration.
Appendix B Additional related works
Personalized FL: Personalized FL has received much attention. In addition to the clustering and local fine-tuning methods introduced in the main paper, we also review other types of personalization techniques used in FL framework.
Model regularization: Hanzely & Richtárik (2020); Sahu et al. (2018); Li et al. (2021) add a regularization term with mixing parameters to penalize the distance between the local and global models. In particular, Sahu et al. (2018) has a pre-set regularization parameter and allows for system heterogeneity, where each client can have a different amount of work to do. Li et al. (2021) consider improving accuracy, while being robust to data and model poisoning attacks, and fair. Similarly, T Dinh et al. (2020) formulate a bi-level optimization problem, which helps decouple personalized model optimization from learning the global model. Huang et al. (2021) propose the FedAMP algorithm which also introduce an additional regularization term, but differ from the previous works in that they encourage similar clients to collaborate more.
Model interpolation: Hanzely & Richtárik (2020) also study a mixed model (local and global model) with a tuning parameter. In their model, as the mixing parameter decreases, it relaxes the local model to be similar to the global model, which can be more personalized. Mansour et al. (2020) propose an idea to combine the global and local model with weight $\alpha$, and Deng et al. (2020) adaptively find the optimal $\alpha^{*}$ as a trade-off at each round for the best performance. Zec et al. (2020); Peterson et al. (2019) both consider using a gating model as a mixing parameter between local and global models. However, Peterson et al. (2019) consider a linear gating model and differentially private FL under domain adaptation, while Zec et al. (2020) split data into two parts used for local and global learning, and they further consider a dropout scenario and the same gating model structure as local and global models.
Data interpolation: As also suggested in Mansour et al. (2020), in addition to the model interpolation, it is possible to combine the local and global data and train a model on their combination. Zhao et al. (2018) create a subset of data that is globally shared across all clients. However, this method is facing the risk of information leaking.
FL with Fairness. In addition to works considering social (group) fairness in FL reviewed in the main paper, we review other types of FL fairness in detail below.
Performance fairness: This line of work measures fairness based on how well the learned model(s) can achieve uniform accuracy across all clients. Li et al. (2019a) propose the $q$-fair FL algorithm which minimizes the aggregate reweighted loss. The idea is that the clients with higher loss will be assigned a higher weight so as to encourage more uniform accuracy across clients. Li et al. (2021) further extend this by considering robustness and poisoning attacks; here, performance fairness and robustness are achieved through a personalized FL method. Zhang et al. (2021) aim to achieve small disparity in accuracy across the groups of client-wise, attribute-wise, and potential clients with agnostic distribution, simultaneously. Wang et al. (2021) discuss the (performance) unfairness caused by conflicting gradients. They detect this conflict through the notion of cosine similarity, and iteratively eliminate it before aggregation by modifying the direction and magnitude of the gradients.
Good-Intent fairness: The good-intent fairness aims to minimize the maximum loss for the protected group. Mohri et al. (2019) propose a new framework of agnostic FL to mitigate the bias in the training procedure via minimax optimization. Similarly, Cui et al. (2021) consider a constrained multi-objective optimization problem to enforce the fairness constraint on all clients. They then maximize the worst client with fairness constraints through a gradient-based procedure. Papadaki et al. (2021) show that a model that is minimax fair w.r.t. clients is equivalent to a relaxed minimax fair model w.r.t. demographic group. They also show their proposed algorithm leads to the same minimax group fairness performance guarantee as the centralized approaches.
Other types of fairness: There are also other types of fairness considered in the FL literature. For instance, Huang et al. (2020) studied the unfairness caused by the heterogeneous nature of FL, which leads to the possibility of preference for certain clients in the training process. They propose an optimization algorithm combined with a double momentum gradient and weighting strategy to create a fairer and more accurate model. Chu et al. (2021) measure fairness as the absolute loss difference between protected groups and labels, a variant of equality opportunity fairness constraint. They propose an estimation method to accurately measure fairness without violating data privacy and incorporate fairness as a constraint to achieve a fairer model with high accuracy performance. Similarly, Zhang et al. (2022) study a new notion of fairness, proportional fairness, in FL, which is based on the relative change of each client’s performance. They connect with the Nash bargaining solution in the cooperative gaming theory and maximize the product of client utilities, where the total relative utility cannot be improved. Similarly, Lyu et al. (2020) study collaborative fairness, meaning that a client who has a higher contribution to learning should be rewarded with a better-performing local model. They introduce a collaborative fair FL framework that incorporates with reputation mechanism to enforce clients with different contributions converge to different models. Their approach could also be viewed as a variant of clustering that separates clients based on their contributions.
Appendix C Dataset and models
In this section, we detail the data and model used in our experiments.
Retiring Adult dataset. We use the pre-processed dataset provided by the folktables Python package (Ding et al., 2021), which provides access to datasets derived from the US Census. In this package, there are three tasks: ACSEmployment, ACSIncome, and ACSHealth. For the ACSEmployemnt task, the goal is to predict whether the person is employed based on its multi-dimensional features; for the ACSIncome task, the goal is to predict whether the person earns more than \$50,000 annually; and for the ACSHealth task, the goal is to predict whether the person is covered by insurance.
Model. We train a fully connected two-layer neural network model for both tasks, where the hidden layer has 32 neurons for the ACSIncome task, and 64 neurons for the ACSEmployment and ACSHealth tasks. For all tasks, we use the RELU activation function and a batch size of 32. Furthermore, we utilize the SGD optimizer for training, with a learning rate of 0.001 for both FedAvg and MAML algorithms and 0.05 for the clustered FL algorithm. In FL, each client updates the global model for 10 epochs in the FedAvg and MAML algorithms and sends it back to the server, while the clustered FL algorithm that has a larger learning rate updates the global model for 1 epoch. We also follow the encoding procedure for categorical features provided by the folktables Python package. The input feature size is 54, 109 and 154 for the ACSIncome, ACSEmployment and ACSHealth tasks, respectively. In the experiments, we consider either sex (e.g., male and female) or race (e.g., White and Non-White) as the protected attribute.
ACSEmployment task with different protected attributes.
Figure 8: Fraction of samples over all states for ACSEmployment
(a) Protected attribute: Race
(b) Protected attribute: Sex
Figure 9: Normalized frequency of fraction of samples for ACSEmployment
As shown in Figure 8 and 9, within the same ACSEmployment task, the data distributions for race (left) and sex (right) are significantly different. For the protected attribute of sex, the number of samples is nearly even across groups and labels. However, for the protected attribute of race, the White group has significantly more samples compared to Non-White groups for both labels ${0,1}$.
ACSIncome and ACSHealth tasks with protected attribute of sex.
We can see from Figure10, 11 and  12 that the fraction of samples in the ACSIncome task is similar across groups for label 0 data but differs significantly for label 1 data. Additionally, we can observe that the ACSHealth task has similar fractions of samples from each group, akin to the ACSEmployment task, in contrast to the ACSIncome task.
(a) Protected attribute: Sex
Figure 10: Normalized frequency of fraction of samples for ACSIncome
(a) Protected attribute: Sex
Figure 11: Normalized frequency of fraction of samples for ACSHealth
Figure 12: Fraction of samples over all states for ACSIncome and ACSHealth
Appendix D Proofs
D.1 Proof of Proposition 1
For simplicity, we assume that clients within the same cluster are identical. Before we show the impact of the global model $\theta^{*}_{G}$ on the fairness performance, we first prove that the global model will lie between two clusters’ models $\theta^{*}_{\alpha}$ and $\theta^{*}_{\beta}$.
Lemma 1.
Under the assumptions of our problem setup, the optimal solution $\theta^{*}_{G}$ for the FedAvg algorithm will lie between $\theta^{*}_{\alpha}$ and $\theta^{*}_{\beta}$.
Proof.
We prove this by contradiction. By definition, let $\theta^{*}_{G}:=\arg\min p*\sum_{j\in\mathcal{C}_{\alpha}}\mathcal{L}_{j}(\theta)+(1-p)*\sum_{j\in\mathcal{C}_{\beta}}\mathcal{L}_{j}(\theta)$ and $\theta^{*}_{i}:=\arg\min\sum_{j\in\mathcal{C}_{i}}\mathcal{L}_{j}(\theta)$ are the optimal solutions for the FedAvg and the clustered FL algorithms, respectively, where $\mathcal{L}_{j}$ is the objective function in Eq. 1. Without loss of generality, we assume $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$. The following proof considers the scenario with $\theta^{*}_{G}>\theta^{*}_{\beta}$; the other case can be shown similarly.
First, it is easy to verify that the objective function is convex in $\theta$. Then, if $\theta^{*}_{G}>\theta^{*}_{\beta}$, it should be that $\sum_{j\in\mathcal{C}_{\beta}}\mathcal{L}_{j}(\theta^{*}_{G})>\sum_{j\in\mathcal{C}_{\beta}}\mathcal{L}_{j}(\theta^{*}_{\beta})$ because $\theta^{*}_{\beta}$ can yield a smaller loss compared to the $\theta^{*}_{G}$. Similarly, we have $\sum_{j\in\mathcal{C}_{\alpha}}\mathcal{L}_{j}(\theta^{*}_{G})>\sum_{j\in\mathcal{C}_{\alpha}}\mathcal{L}_{j}(\theta^{*}_{\beta})>\sum_{j\in\mathcal{C}_{\alpha}}\mathcal{L}_{j}(\theta^{*}_{\alpha})$ due to convexity. Therefore, $\theta^{*}_{G}$ is not the optimal solution, contradicting the assumption. Hence, the FedAvg solution would lie between $\theta^{*}_{\alpha}$ and $\theta^{*}_{\beta}$. ∎
Now, we are ready to prove the Proposition 1 that global model is more unfair than the cluster-specific model (i.e., $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})<\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$).
Proof.
We start with the scenario where $r_{a}=r_{b}$, balanced label participation rates, and equalized distance between peaks. As the following analysis focuses on the cluster $\mathcal{C}_{\alpha}$, we drop the cluster notation from the derivation for notation simplicity. Let $\Delta_{\texttt{EqOp}}(\theta)$ be the cluster-wise EqOp fairness gap at the given decision threshold $\theta$. Based on the definition, it could be written as
$\Delta_{\texttt{EqOp}}(\theta)=\int_{\theta}^{\infty}f^{1}_{a}(x)\mathrm{d}x-\int_{\theta}^{\infty}f^{1}_{b}(x)\mathrm{d}x.$
According to the Leibniz integral rule, we can find the derivative of $\Delta_{\texttt{EqOp}}(\theta)$ w.r.t. $\theta$ as following:
$\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta)=f^{1}_{b}(\theta)-f^{1}_{a}(\theta)$
Let the intersection point of the feature-label distribution $f^{y}_{g}$ and $f^{y^{\prime}}_{g^{\prime}}$ be $I_{g^{y},{g^{\prime}}^{y^{\prime}}}$. It is easy to verify that the optimal decision threshold $\theta^{*}_{\alpha}$ obtained from  1 could be written in the closed form such that
$\theta^{*}_{\alpha}=I_{a^{1},{b}^{0}}=I_{b^{1},{a}^{0}}$
When $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta)=0$, $\theta=\infty,-\infty$ or $I_{a^{1},{b}^{1}}$. Furthermore, at extreme cases where $\theta\rightarrow\infty$ or $\infty$- ∞, we can find that the value of EqOp fairness gap $\Delta_{\texttt{EqOp}}(\infty)=\Delta_{\texttt{EqOp}}(-\infty)=0$. Therefore, to investigate the impact of FedAvg solution $\theta^{*}_{G}$ on the EqOp fairness gap, it is equivalent to check the sign of $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta)$ at the optimal decision threshold $\theta^{*}_{\alpha}$ obtained by solving  1.
To relax the equalized distance assumption, we could treat the location of modes of $f^{1}_{a},f^{0}_{a},f^{1}_{b}$ as fixed, and vary the mode of $f^{0}_{b}$, and there are two cases we can discuss:
  1. 1.
    $I_{a^{1},{b}^{1}}-I_{a^{0},{b}^{1}}<I_{a^{0},{b}^{1}}-I_{a^{0},{b}^{0}}$
    Under this condition, we could consider a smaller value of the mode of $f^{0}_{b}$. As a result, the optimal decision threshold $\theta^{*}_{\alpha}$ will shift to the left, resulting in a smaller value compared to the equalized distance case. In other words, it means $\theta^{*}_{\alpha}\leq I_{a^{1},{b}^{1}}$, indicating $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta^{*}_{\alpha})\geq 0$.
  2. 2.
    $I_{a^{1},{b}^{1}}-I_{a^{0},{b}^{1}}>I_{a^{0},{b}^{1}}-I_{a^{0},{b}^{0}}$
    Under this condition, we could consider a larger value of the mode of $f^{0}_{b}$, but it is still less than that of $f^{0}_{a}$ according to our assumption. At the extreme case when they are equal, the optimal decision threshold determined from  1 would be smaller than $I_{a^{0},{a}^{1}}$ because of the mode of $f^{1}_{b}$ is less than that of $f^{1}_{a}$, which is also smaller than $I_{a^{1},{b}^{1}}$. Therefore, we can still conclude $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta^{*}_{\alpha})\geq 0$.
For the scenario of $r_{a}\neq r_{b}$, we can find that the change of $r_{g}$ does not affect the value of $\Delta_{\texttt{EqOp}}(\theta)$, but the location of $\theta^{*}_{\alpha}$. According to our distribution assumption, when $r_{a}\geq(\text{resp.}\leq)r_{b}$, the optimal solution $\theta^{*}_{\alpha}$ will be in favor of the group $a$ (resp. $b$) distributions, leading to a right (resp. left) shift compared to the optimal solution when $r_{a}=r_{b}$. However, when $r_{a}\rightarrow$ 1 (resp. 0), $\theta^{*}_{\alpha}\rightarrow I_{a^{0},{a}^{1}}$ (resp. $I_{b^{0},{b}^{1}}$), which is still less than $I_{a^{1},{b}^{1}}$, indicating $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta^{*}_{\alpha})\geq 0$.
With the assumption that the majority of samples are labeled as 1 (i.e., $\alpha^{1}_{g}\geq\alpha^{0}_{g}$), the decision threshold $\theta^{*}_{\alpha}$ will shift towards the left to account for label imbalance. In other words, the sign of $\Delta_{\texttt{EqOp}}^{{}^{\prime}}(\theta^{*}_{\alpha})$ remains positive. Since $\theta^{*}_{\alpha}<\theta^{*}_{\beta}$, there exist a cluster size weight $p$ such that the FedAvg solution $\theta^{*}_{G}$ will make the cluster $\mathcal{C}_{\alpha}$ unfairer. ∎
D.2 Proof of Proposition 2
Proof.
For simplicity, we assume that clients within the same cluster are identical. We start with the scenario where $r_{a}=r_{b}$ and balanced label participation rate. As the following analysis focuses on the cluster $\mathcal{C}_{\alpha}$, we drop the cluster notation from the derivation for notation simplicity. Let $\Delta_{\texttt{SP}}(\theta)$ be the cluster-wise SP fairness gap at the given decision threshold $\theta$. According to its definition, it could be written as
$\Delta_{\texttt{SP}}(\theta)=\alpha^{1}_{a}\int_{\theta}^{\infty}f^{1}_{a}(x)\mathrm{d}x+\alpha^{0}_{a}\int_{\theta}^{\infty}f^{0}_{a}(x)\mathrm{d}x-\alpha^{1}_{b}\int_{\theta}^{\infty}f^{1}_{b}(x)\mathrm{d}x-\alpha^{0}_{b}\int_{\theta}^{\infty}f^{0}_{b}(x)\mathrm{d}x.$
According to the Leibniz integral rule, we can find the derivative of $\Delta_{\texttt{SP}}(\theta)$ w.r.t. $\theta$ as following:
$\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta)=\alpha^{1}_{b}f^{1}_{b}(\theta)+\alpha^{0}_{b}f^{0}_{b}(\theta)-\alpha^{1}_{a}f^{1}_{a}(\theta)-\alpha^{0}_{a}f^{0}_{a}(\theta)$
According to our distribution assumptions, we can write the above expression in the following closed form with $\alpha=\alpha^{y}_{g}\hskip 3.61371pt\forall y,g$
$\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta)=\frac{\alpha}{\sqrt{2\pi}\sigma}\Big{(}\exp(-\frac{(\theta-\mu^{1}_{b})^{2}}{2\sigma^{2}})+\exp(-\frac{(\theta-\mu^{0}_{b})^{2}}{2\sigma^{2}})-\exp(-\frac{(\theta-\mu^{1}_{a})^{2}}{2\sigma^{2}})-\exp(-\frac{(\theta-\mu^{0}_{a})^{2}}{2\sigma^{2}})\Big{)}$
Furthermore, it is easy to verify that the optimal decision threshold $\theta^{*}_{\alpha}$ obtained by solving  1 could be written in the closed form such that
$\bar{\theta}=\theta^{*}_{\alpha}=\frac{\mu^{1}_{a}+\mu^{0}_{b}}{2}=\frac{\mu^{1}_{b}+\mu^{0}_{a}}{2}$
At the optimal solution $\theta^{*}_{\alpha}$, $\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta^{*}_{\alpha})=0$. Similar to the proof of Proposition 1, to investigate the impact of FedAvg solution $\theta^{*}_{G}$ on the SP fairness gap, it is equivalent to check how the $\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta^{*}_{\alpha})$ change in the neighborhood of the optimal solution $\theta^{*}_{\alpha}$. Also, at extreme cases, we can easily find that the value of SP fairness gap $\Delta_{\texttt{SP}}(\infty)=\Delta_{\texttt{SP}}(-\infty)=0$. Therefore, if $\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta^{*}_{\alpha})\geq 0$, then we can conclude that the FedAvg solution $\theta^{*}_{G}$ would lead to a worse fairness performance compared to the optimal solution $\theta^{*}_{\alpha}$. Let $\psi_{1}(\theta)=\exp(-\frac{(\theta-\mu^{1}_{b})^{2}}{2\sigma^{2}})-\exp(-\frac{(\theta-\mu^{0}_{a})^{2}}{2\sigma^{2}})$ and $\psi_{2}(\theta)=\exp(-\frac{(\theta-\mu^{1}_{a})^{2}}{2\sigma^{2}})-\exp(-\frac{(\theta-\mu^{0}_{b})^{2}}{2\sigma^{2}})$. At the solution $\theta^{*}_{\alpha}$, we can find that $\psi_{1}(\theta^{*}_{\alpha})=\psi_{2}(\theta^{*}_{\alpha})=0$.
Hence, to investigate how the $\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta^{*}_{\alpha})$ change, we can find the rate of change for both $\psi_{1}(\theta)$ and $\psi_{2}(\theta)$ in the neighborhood of $\theta^{*}_{\alpha}$ such that
$\psi^{{}^{\prime}}_{1}(\theta^{*}_{\alpha})=\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{0}_{a})^{2}}{-2\sigma^{2}})\tfrac{\theta^{*}_{\alpha}-\mu^{0}_{a}}{\sigma}-\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{1}_{b})^{2}}{-2\sigma^{2}})\tfrac{\theta^{*}_{\alpha}-\mu^{1}_{b}}{\sigma}=\tfrac{1}{\sigma}\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{1}_{b})^{2}}{-2\sigma^{2}})(\mu^{1}_{b}-\mu^{0}_{a})$
$\psi^{{}^{\prime}}_{2}(\theta^{*}_{\alpha})=\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{0}_{b})^{2}}{-2\sigma^{2}})\tfrac{\theta^{*}_{\alpha}-\mu^{0}_{b}}{\sigma}-\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{1}_{a})^{2}}{-2\sigma^{2}})\tfrac{\theta^{*}_{\alpha}-\mu^{1}_{a}}{\sigma}=\tfrac{1}{\sigma}\exp(\tfrac{(\theta^{*}_{\alpha}-\mu^{0}_{b})^{2}}{-2\sigma^{2}})(\mu^{1}_{a}-\mu^{0}_{b})$
By setting $\psi^{{}^{\prime}}_{1}(\theta^{*}_{\alpha})\geq\psi^{{}^{\prime}}_{2}(\theta^{*}_{\alpha})$, it means the increment of $\psi_{1}$ is larger than the decrement of $\psi_{2}$. Therefore, with Lemma 1, there exists a cluster size weight $p$ such that the FedAvg solution $\theta^{*}_{G}$ will make the cluster $\mathcal{C}_{\alpha}$ unfairer. The inequality is obtained by considering unequalized $\alpha^{y}_{g}$.
In addition, for the scenario of $r_{a}\neq r_{b}$, similar to the proof of Proposition 1, we can find that the change of $r_{g}$ does not affect the expression of $\Delta_{\texttt{SP}}(\theta)$, but it will affect the location of $\theta^{*}_{\alpha}$. According to our distribution assumption, when $r_{a}\geq(\text{resp.}\leq)r_{b}$, the optimal solution $\theta^{*}_{\alpha}$ will be in favor of the group $a$ (resp. $b$) distributions, leading to a right (resp. left) shift compared to the optimal solution when $r_{a}=r_{b}$. However, when $r_{a}\rightarrow$ 1 (resp. 0), $\theta^{*}_{\alpha}\rightarrow\frac{\mu^{0}_{a}+\mu^{1}_{a}}{2}$ (resp. $\frac{\mu^{0}_{b}+\mu^{1}_{b}}{2}$), which is limited within the range of $(\frac{\mu^{0}_{a}+\mu^{0}_{b}}{2},\frac{\mu^{1}_{a}+\mu^{1}_{b}}{2})$. When $\theta=\frac{\mu^{1}_{a}+\mu^{1}_{b}}{2}$, we can easily find that $\Delta_{\texttt{SP}}^{{}^{\prime}}(\theta)\approx 0$ especially when $\sigma$ is small. In other words, we can conclude that $\Delta_{\texttt{SP}}(\theta)\geq\Delta_{\texttt{SP}}(\theta^{*}_{\alpha})$ for any $\theta^{*}_{\alpha}\in(\frac{\mu^{0}_{b}+\mu^{1}_{b}}{2},\frac{\mu^{0}_{a}+\mu^{1}_{a}}{2})$. Therefore, the claim still holds.
Furthermore, when the equalized label participation rate assumption is relaxed, the above proof strategy still holds by considering different $\alpha^{y}_{g}$ into the expression. It is worth noting that when the label participation rates are balanced, the fairness $\Delta_{\texttt{SP}}(\theta)$ has two equal-height peaks (e.g., $\Delta_{\texttt{SP}}^{\prime}(\theta)=0$) by symmetricity of the Gaussian distribution when $\theta\approx\frac{\mu^{1}_{a}+\mu^{1}_{b}}{2}$ and $\frac{\mu^{0}_{a}+\mu^{0}_{b}}{2}$. However, when the majority of samples are labeled as 1, we observe a shift in the decision threshold $\theta^{*}_{\alpha}\leq\bar{\theta}$ towards the left to account for label imbalance. In this case, since $\theta^{*}_{G}>\theta^{*}_{\alpha}$, the FedAvg solution pulls $\theta^{*}_{\alpha}$ upwards, favoring label 1, which results in both accuracy and fairness deteriorating. Moreover, when $r_{b}\geq r_{a}$ and the majority of samples are labeled 1 in one group where the other group has a better balance in the label (i.e., $\alpha^{1}_{a}>\alpha^{0}_{b}>\alpha^{1}_{b}>\alpha^{0}_{a}$), $\theta^{*}_{\alpha}\leq\bar{\theta}$ holds. Therefore, $\Delta_{\texttt{SP}}(\theta)$ will increase initially and then decrease, and there still exist a cluster size weight $p$ such that the FedAvg solution $\theta^{*}_{G}$ will make the cluster $\mathcal{C}_{\alpha}$ unfairer. ∎
Appendix E Additional numerical experiments
E.1 Additional experiments on other types of fairness notions
(a) Employment-sex: EO
(b) Income-sex: EO
(c) Employment-sex: EqOp
Figure 13: Personalization could also improve other fairness notions
In addition to the SP fairness investigated in Section 4, we also study the impact of personalization techniques on other types of fairness notions such as EO and EqOp. From Fig. 13, we find that the introduction of personalization techniques can enhance other types of fairness due to the computational advantages of collaboration. However, compared to the improvement of SP and EqOp fairness, the local EO fairness improvement is less significant because the EO matches both the true and false positive rates across two protected groups, rendering it a more stringent criterion.
E.2 Additional experiments on other datasets
(a) Health-sex: SP
(b) Income-race: SP
Figure 14: Additional experiments on other datasets with different protected attributes
In addition to the ACSEmployment (sex, race) and ACSIncome (sex) experimental results presented in Section 4, we conducted additional experiments to explore the impact of SP fairness using new datasets, as illustrated in Fig. 14. Examining the ACSHealth data with sex as the protected attribute, we can see from Fig. 12 that the fractions of samples from each group across all states are similar to that of the ACSEmployment dataset shown in Fig. 8, resulting in a similar performance. From Fig. 14, we can see that personalization techniques can improve local fairness as an unintended benefit, similar to the observations from Section 4.
E.3 Additional experiments on the EqOp and EO fairness
In Section 6, we compare SP fairness between two algorithms: ICFA and Fair-FCA. Here, we also compare the EqOp and EO fairness between them. The observations from Table 1 are also consistent with those from Fig 7, meaning that the Fair-FCA algorithm enables us to establish a better fairness-accuracy tradeoff (a drop in accuracy in return for improved fairness) compared to the IFCA algorithm.
Table 1: Fair-FCA with different datasets and protected attributes
Dataset Algorithm EqOp Acc. (EqOp) EO Acc. (EO)
Employment-Race IFCA 0.07764 0.8188 0.09319 0.8188
Fair-FCA 0.07029 0.8151 0.08946 0.8151
Employment-Sex IFCA 0.04521 0.8188 0.05808 0.8188
Fair-FCA 0.04183 0.8157 0.05655 0.8151
Income-Sex IFCA 0.05029 0.7511 0.05231 0.7511
Fair-FCA 0.04932 0.7491 0.05161 0.7489
E.4 Details of setup on synthetic experiment
Table 2: Data distributions over 8 clients
Client ID $f^{1}_{1}$ $f^{0}_{1}$ $f^{1}_{0}$ $f^{0}_{0}$
1 $N(8,1)$ $N(6,1)$ $N(8,1)$ $N(6,1)$
2 $N(12,1)$ $N(8,1)$ $N(11,1)$ $N(7,1)$
3 $N(7.5,1)$ $N(5.5,1)$ $N(7.5,1)$ $N(5.5,1)$
4 $N(12,1)$ $N(9,1)$ $N(12,1)$ $N(9,1)$
5 $N(12,1)$ $N(8,1)$ $N(11,1)$ $N(7,1)$
6 $N(11.5,1)$ $N(8.5,1)$ $N(11.5,1)$ $N(8.5,1)$
7 $N(11,1)$ $N(8,1)$ $N(11,1)$ $N(8,1)$
8 $N(10.5,1)$ $N(7.5,1)$ $N(10.5,1)$ $N(7.5,1)$
According to the data distribution information, we can see that clients 2,4,5,6,7,8 have similar data distributions compared to clients 1,3. Also, we can find that clients 1,3,4,6,7,8 share identical data distribution across the two groups. We generate 1200 samples from each distribution and apply a logistic regression classifier for binary classification tasks. We report our experiment results for an average of 5 runs. When $\gamma=1$, Fair-FCA prioritizes accuracy; by design, this is attained by grouping the 6 clients having similar data distributions together ({1,3} and {2,4,5,6,7,8}). Similarly, when $\gamma=0$, Fair-FCA focuses only on SP fairness, this time clustering clients that have identical distributions on the two protected groups together ({2,5} and {1,3,4,6,7,8}). Lastly, by setting $\gamma\in(0,1)$, we can effectively account for both accuracy and SP fairness when clustering: when $\gamma=0.3$, the clusters are {2,4,5} and {1,3,6,7,8}; when $\gamma=0.5$, the clusters are {2,4,5,6} and {1,3,7,8}; and when $\gamma=0.8$, the clusters are {2,4,5,6,7} and {1,3,8}.
Figure 15: Larger version of Figure 6; Accuracy vs. SP-Fairness
E.5 Additional experiments using original Retiring Adult dataset without feature scaling
Table 3: Algorithm performance comparisons using original Retiring adult dataset
Dataset Algorithm SP Acc. (SP) EqOp Acc. (EqOp)
Employment-Sex IFCA 0.03667 0.8229 0.04698 0.8229
Fair-FCA 0.03594 $\downarrow$ 0.8223 $\downarrow$ 0.04633 $\downarrow$ 0.8224 $\downarrow$
Employment-Race IFCA 0.07257 0.8229 0.07315 0.8229
Fair-FCA 0.07219 $\downarrow$ 0.8224 $\downarrow$ 0.06527 $\downarrow$ 0.8226 $\downarrow$
Income-Sex IFCA 0.08355 0.7481 0.04773 0.7481
Fair-FCA 0.08227 $\downarrow$ 0.7469 $\downarrow$ 0.04767 $\downarrow$ 0.7469 $\downarrow$
Income-Race IFCA 0.1012 0.7481 0.1100 0.7481
Fair-FCA 0.1011 $\downarrow$ 0.7468 $\downarrow$ 0.1086 $\downarrow$ 0.7466 $\downarrow$
We can see from Table 3 that compared to the IFCA algorithm, our Fair-FCA algorithm is experiencing a degradation in accuracy but an improved fairness, meaning an accuracy-fairness tradeoff. These observations are also consistent with our findings when using the Retiring Adult dataset with feature scaling in Section 6.
E.6 Additional discussions on the assumptions used for the analytical results
(a) SP
(b) EqOp
Figure 16: Fairness under different values of decision threshold
In Fig.16, both cluster-wise SP and EqOp fairness and loss are assessed under the different values of decision threshold $\theta$ on local data, depicted in red and green colors. The solid lines denote the corresponding fairness performance, while the dotted lines illustrate the corresponding loss under the corresponding decision threshold. The vertical red/green lines show the optimal decision threshold determined by solving Eq. 1, with the blue vertical line indicating the FedAvg solution with the parameter $p=0.5$.
For a more insightful comparison, we assume that in the cluster $\mathcal{C}_{\beta}$ (green), the data distribution exhibits an equalized distance between distributions, equalized group rate, and equalized label rate (e.g., $(\mu^{1}_{a},\mu^{1}_{b},\,\mu^{0}_{a},\mu^{0}_{b})=(10,9,7,6),\alpha^{y}_{g}=0.5,r_{g}=0.5$). Meanwhile, in the cluster $\mathcal{C}_{\alpha}$ (red), we relax all of these assumptions (e.g., $(\mu^{1}_{a},\mu^{1}_{b},\,\mu^{0}_{a},\mu^{0}_{b})=(7,6,4,2),(\alpha^{1}_{a},\alpha^{1}_{b},\alpha^{0}_{a},\alpha^{0}_{b})=(0.6,0.3,0.4,0.7),(r_{a},r_{b})=(0.65,0.35)$). The comparison depicted in Fig. 16 reveals that the SP fairness exhibits less structured changes as the decision moves (e.g., due to the used of a global model). This is because EqOp fairness solely considers the true positive rates across two protected groups. The observation also underscores the necessity for more restrictive assumptions in our analytical support for SP fairness (Proposition 2).
E.7 Experiments under Gaussian distribution with equalized distance, balanced group rate, and equalized label rate
Numerical illustration. We now conduct numerical experiments to illustrate the findings in Prop. 12. We drop the cluster notation $c$ whenever it is clear from the context. The results are presented in Tables 4 and 5. We proceed as follows: 10000 random samples in cluster $\mathcal{C}_{\alpha}$ are drawn from Gaussian distribution for each group $g\in\{a,b\}$ with mean $\mu^{y,\mathcal{C}_{\alpha}}_{g}$ and standard deviation $\sigma$. The number of qualified ($y=1$) and unqualified ($y=0$) samples in each group is proportional to the label participation rate $\alpha^{y,\mathcal{C}_{\alpha}}_{g}$. Since samples were generated in a consistent manner across different parameter settings, we assumed an optimal decision threshold $\theta^{*}_{\beta}=8$ for cluster $\mathcal{C}_{\beta}$, obtained according to the distribution information: $(f^{1}_{1},f^{0}_{1},f^{1}_{0},f^{0}_{0},\sigma)=(10,7,9,6,1)$ with equalized group rate $r_{g}=0.5,\forall g$ and label participation rate $\alpha^{y,\mathcal{C}_{\beta}}_{g}=0.5,\forall g,y$. In Table 4, we consider the scenario where $\alpha^{y,\mathcal{C}_{\alpha}}_{g}=0.5\hskip 3.61371pt\forall g,y$. In contrast, different values of $\alpha^{y,\mathcal{C}_{\alpha}}_{g}$ are applied in Table 5. Both results in Table 4 and  5 consider an equalized group rate such that $r_{a}=r_{b}$ and an equalized distance between mean estimates.
From Table 4, we can find that it offers crucial insights into the conditions required for Proposition 2 (SP) to hold. For fixed mean estimates $\mu^{y}_{g}$ (rows 1-2), we observe that smaller values of $\sigma$ are preferable to satisfy the specified conditions. Similarly, for fixed $\sigma$ (row 1, 3 and row 2, 4), larger differences between $\mu^{1}_{g}$ and $\mu^{0}_{g}$ are advantageous in fulfilling the conditions. This observation becomes intuitive at the extreme cases where samples are linearly separable with small $\sigma$ or large distance between $\mu^{1}_{g}$ and $\mu^{0}_{g}$. Therefore, the optimal decision threshold $\theta^{*}_{\alpha}$ could achieve a perfect classification as well as perfect fairness. Hence, the FedAvg solution $\theta^{*}_{G}$ deviated from the optimal solution will lead to worse performance in both accuracy and fairness. We could also observe that for the EqOp fairness, under an equalized label rate, the FedAvg solutions consistently make the cluster $\mathcal{C}_{\alpha}$ unfairer, which is consistent with our findings in Prop. 1.
Table 4: Cluster $\mathcal{C}_{\alpha}$ fairness performance with equalized distance, group rate and label rate
Distribution Condition $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) (SP) $p=\frac{2}{3}$ $p=\frac{1}{2}$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4, 6, 3, 1) Yes 0.1359 0.1814 $\uparrow$ 0.1945 $\uparrow$ 0.1359 0.3413 $\uparrow$ 0.3829 $\uparrow$
(7, 4, 6, 3, 2) No 0.1499 0.1417 $\downarrow$ 0.1315 $\downarrow$ 0.1499 0.1915 $\uparrow$ 0.1974 $\uparrow$
(7, 5, 6, 4, 1) No 0.2417 0.2297 $\downarrow$ 0.2046 $\downarrow$ 0.2417 0.3781 $\uparrow$ 0.3721 $\uparrow$
(8, 3, 6, 1, 2) Yes 0.1866 0.1968 $\uparrow$ 0.2033 $\uparrow$ 0.1866 0.3121 $\uparrow$ 0.3590 $\uparrow$
Table 5 reveals insights regarding the influence of label distribution $\alpha^{y}_{g}$ on SP and EqOp fairness performance. Specifically, when the majority of samples in both groups are labeled as 1 (rows 1-2), the optimal decision threshold ($\theta^{*}_{\alpha}$) shifts leftward compared to the balanced scenario. However, with Lemma 1, the FedAvg solution $\theta^{*}_{G}$ will be greater than $\theta^{*}_{\alpha}$. Therefore, we can find that $\theta$ will have even larger fairness gap when it is shifted to the right. Another intriguing observation is that in cases where the majority of samples have different labels (row 3), the FedAvg solution ($\theta^{*}_{G}$) yields worse fairness performance when $p=2/3$ or $1/2$ but not when $p=1/3$ (0.1720 $\downarrow$) or $1/4$ (0.1391 $\downarrow$). This indicates the weight $p$ plays a significant role in shaping the overall cluster-wise average fairness performance, especially when assessing the overall cluster-wise average fairness performance.
Table 5: Cluster $\mathcal{C}_{\alpha}$ fairness performance with equalized distance and group rate
Distribution Label rate Condition $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) (SP) $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4, 6, 3, 1) (0.7, 0.3, 0.6, 0.4) Yes 0.2062 0.2146 $\uparrow$ 0.2463 $\uparrow$
(0.6, 0.4, 0.7, 0.3) Yes 0.0453 0.0514 $\uparrow$ 0.0813 $\uparrow$
(0.7, 0.3, 0.4, 0.6) Yes 0.3797 0.3858 $\uparrow$ 0.3926 $\uparrow$
(0.6, 0.4, 0.3, 0.7) No 0.3797 0.3748 $\downarrow$ 0.2804 $\downarrow$
(EqOp) $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$
(7, 4, 6, 3, 2) (0.7, 0.3, 0.6, 0.4) Yes 0.0998 0.1807 $\uparrow$ 0.1923 $\uparrow$
(0.6, 0.4, 0.7, 0.3) Yes 0.0975 0.1198 $\uparrow$ 0.1796 $\uparrow$
(0.1, 0.9, 0.5, 0.5) No 0.1965 0.1650 $\downarrow$ 0.1574 $\downarrow$
(0.3, 0.7, 0.2, 0.8) No 0.1974 0.1645 $\downarrow$ 0.1574 $\downarrow$
Since we assume clients within the same cluster are identical, and the local fairness performance for an algorithm can be computed as a weighted sum of the local fairness performance from each cluster, the cluster-wise average local fairness gap under different models’ optimal solution $\theta$ could be calculated as $\Delta_{f}(\theta)=p\Delta^{\alpha}_{f}+(1-p)\Delta^{\beta}_{f};f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$, where $p$ is the fraction of clients belonging to cluster $\mathcal{C}_{\alpha}$.
In Table 6 and  7, we delve into different notions of cluster-wise average fairness gap achieved with different decision thresholds (optimal clustered FL solutions $\theta^{*}_{C}$ and FedAvg solutions $\theta^{*}_{G}$). In the following experiment, we keep the parameters in cluster $\mathcal{C}_{\beta}$ as constants while varying those in cluster $\mathcal{C}_{\alpha}$ to assess its impact on the corresponding fairness. From the results in Table 6 and  7, we can find that when both conditions are not satisfied (rows 5-6), there is a cluster size weight $p$ such that the FedAvg solutions would lead to better fairness performance for each cluster, consequently yielding a lower cluster-wise average fairness gap. However, when only one cluster satisfies the condition, meaning that there is a $p$ such that the FedAvg solutions would only make one cluster unfairer (rows 1-2 in Table 6), we could see that a relatively small $p$ would let the clustered FL solutions yield a better fairness performance because $\theta^{*}_{G}$ will move to the cluster with a smaller value of $p$ to account for the cluster size imbalance. Nevertheless, when $p$ is large, the FedAvg solutions will again have superior fairness performance than the clustered FL solutions, similar to the results in rows 3-4 in Table 6 and  7. Essentially, for each cluster $c$, there exists a range $(p^{c}_{low},p^{c}_{high})$ such that, within this range, FedAvg solutions result in worse fairness performance compared to clustered FL solutions. Consequently, for any $p\in\cap_{c}(p^{c}_{low},p^{c}_{high})$, clustered FL solutions yield a superior cluster-wise average fairness performance relative to FedAvg solutions.
Table 6: Cluster-wise average SP fairness performance with equalized distance
Distribution Label rate Condition $p$ $\Delta_{\texttt{SP}}(\theta^{*}_{C})$ $\Delta_{\texttt{SP}}(\theta^{*}_{G})$
$\mathcal{C}_{\alpha}:$($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$)
$\mathcal{C}_{\beta}:$($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) satisfied
(7, 4, 6, 3, 2) (0.5, 0.5, 0.5, 0.5) No 4/5 0.147 0.145 $\downarrow$
(10, 7, 9, 6, 1) (0.5, 0.5, 0.5, 0.5) Yes 1/3 0.141 0.160 $\uparrow$
(7, 4, 6, 3, 2) (0.8, 0.2, 0.7, 0.3) Yes 3/4 0.139 0.107 $\downarrow$
(10, 7, 9, 6, 1) (0.5, 0.5, 0.5, 0.5) Yes 1/2 0.138 0.178 $\uparrow$
(7, 4, 6, 3, 2) (0.5, 0.5, 0.5, 0.5) No 1/3 0.303 0.283 $\downarrow$
(10, 7, 9, 6, 1) (0.7, 0.3, 0.4, 0.6) No 2/3 0.227 0.200 $\downarrow$
Table 7: Cluster-wise average EqOp fairness performance with equalized distance
Distribution Label rate Condition $p$ $\Delta_{\texttt{EqOp}}(\theta^{*}_{C})$ $\Delta_{\texttt{EqOp}}(\theta^{*}_{G})$
$\mathcal{C}_{\alpha}:$($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$)
$\mathcal{C}_{\beta}:$($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) satisfied
(7, 4, 6, 3, 2) (0.3, 0.7, 0.2, 0.8) No 1/3 0.156 0.133 $\downarrow$
(10, 7, 9, 6, 1) (0.5, 0.5, 0.5, 0.5) Yes 2/3 0.177 0.139 $\downarrow$
(7, 4, 6, 3, 2) (0.8, 0.2, 0.7, 0.3) Yes 3/4 0.082 0.050 $\downarrow$
(10, 7, 9, 6, 1) (0.5, 0.5, 0.5, 0.5) Yes 1/2 0.100 0.109 $\uparrow$
(7, 4, 6, 3, 2) (0.3, 0.7, 0.2, 0.8) No 1/3 0.224 0.187 $\downarrow$
(10, 7, 9, 6, 1) (0.3, 0.7, 0.2, 0.8) No 2/3 0.211 0.149 $\downarrow$
E.8 Additional experiments under Gaussian distribution with equalized distance and balanced label rate
Compared to the experiments focused on an all balanced setting in Table 4, the following experiments relax the group rates setting in the cluster $\mathcal{C}_{\alpha}$, while we keep other settings (i.e., balanced label rate and equalized distance) and data information for $\mathcal{C}_{\beta}$ unchanged.
Table 8: Cluster $\mathcal{C}_{\alpha}$ fairness performance under Gaussian distribution with equalized distance and label rate, but not group rate
Distribution Group rate $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) $(r_{a},r_{b})$ $p=\frac{2}{3}$ $p=\frac{1}{2}$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4, 6, 3, 1) (0.5, 0.5) 0.1359 0.1814 $\uparrow$ 0.1945 $\uparrow$ 0.1359 0.3413 $\uparrow$ 0.3829 $\uparrow$
(0.7, 0.3) 0.1388 0.1558 $\uparrow$ 0.1941 $\uparrow$ 0.1780 0.2594 $\uparrow$ 0.3828 $\uparrow$
(0.9, 0.1) 0.1465 0.1702 $\uparrow$ 0.1941 $\uparrow$ 0.2217 0.3076 $\uparrow$ 0.3828 $\uparrow$
(0.3, 0.7) 0.1388 0.1359 $\downarrow$ 0.1558 $\uparrow$ 0.0996 0.1359 $\uparrow$ 0.2594 $\uparrow$
(0.4, 0.6) 0.1367 0.1372 $\uparrow$ 0.1931 $\uparrow$ 0.1161 0.1634 $\uparrow$ 0.3759 $\uparrow$
From Table 8, we can see that the changes in the group rate do not affect the fairness performance comparison. There exists a cluster size weight $p$ such that the FedAvg solutions would lead to worse SP and EqOp fairness performance compared to the clustered FL solutions. This observation is also consistent with our findings in the Proposition 1 and  2.
E.9 Additional experiments under Gaussian distribution with equalized distance
Similar to experiments in E.8, we further relax balanced label rate setting in the following experiments, while we keep other settings (i.e., equalized distance) and data information for $\mathcal{C}_{\beta}$ unchanged.
Table 9: Cluster $\mathcal{C}_{\alpha}$ SP fairness performance under Gaussian distribution with equalized distance, but not label rate and group rate
Distribution Label rate Group rate $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) $(r_{a},r_{b})$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4, 6, 3, 1) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.2062 0.2146 $\uparrow$ 0.2463 $\uparrow$
(0.3, 0.7) 0.2024 0.2056 $\uparrow$ 0.2167 $\uparrow$
(0.7, 0.3) 0.2136 0.2309 $\uparrow$ 0.2793 $\downarrow$
(0.6, 0.4, 0.7, 0.3) (0.5, 0.5) 0.0453 0.0514 $\uparrow$ 0.0813 $\uparrow$
(0.3, 0.7) 0.0460 0.0446 $\downarrow$ 0.0482 $\uparrow$
(0.7, 0.3) 0.0535 0.0751 $\uparrow$ 0.1467 $\uparrow$
(0.7, 0.3, 0.4, 0.6) (0.5, 0.5) 0.3797 0.3858 $\uparrow$ 0.3926 $\uparrow$
(0.3, 0.7) 0.3780 0.3819 $\uparrow$ 0.3451 $\downarrow$
(0.7, 0.3) 0.3821 0.3899 $\uparrow$ 0.3936 $\uparrow$
(0.4, 0.6, 0.7, 0.3) (0.7, 0.3) 0.1005 0.0662 $\downarrow$ 0.0766 $\downarrow$
(0.9, 0.1) 0.0725 0.0078 $\downarrow$ 0.0868 $\uparrow$
(0.3, 0.7) 0.1013 0.1084 $\uparrow$ 0.1090 $\downarrow$
(0.1, 0.9) 0.0767 0.0860 $\uparrow$ 0.0972 $\uparrow$
Table 10: Cluster $\mathcal{C}_{\alpha}$ EqOp fairness performance under Gaussian distribution with equalized distance, but not label rate and group rate
Distribution Label rate Group rate $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) $(r_{a},r_{b})$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4, 6, 3, 2) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.0998 0.1222 $\uparrow$ 0.1807 $\uparrow$
(0.3, 0.7) 0.0952 0.1109 $\uparrow$ 0.1784 $\uparrow$
(0.7, 0.3) 0.1044 0.1386 $\uparrow$ 0.1825 $\uparrow$
(0.6, 0.4, 0.7, 0.3) (0.5, 0.5) 0.0975 0.1198 $\uparrow$ 0.1796 $\uparrow$
(0.3, 0.7) 0.0798 0.0874 $\uparrow$ 0.1799 $\uparrow$
(0.7, 0.3) 0.1180 0.1957 $\uparrow$ 0.1792 $\uparrow$
(0.1, 0.9, 0.5, 0.5) (0.5, 0.5) 0.1965 0.1650 $\downarrow$ 0.1574 $\downarrow$
(0.3, 0.7) 0.1751 0.1742 $\downarrow$ 0.1620 $\downarrow$
(0.7, 0.3) 0.1869 0.1569 $\downarrow$ 0.1537 $\downarrow$
(0.3, 0.7, 0.2, 0.8) (0.5, 0.5) 0.1974 0.1645 $\downarrow$ 0.1574 $\downarrow$
(0.3, 0.7) 0.1974 0.1630 $\downarrow$ 0.1569 $\downarrow$
(0.7, 0.3) 0.1973 0.1660 $\downarrow$ 0.1585 $\downarrow$
From Table 9, we can observe that for the SP fairness, when the majority of samples are labeled 1 (rows 1-6), the changes in the group rate do not affect the fairness performance comparison in the cluster $\mathcal{C}_{\alpha}$. There exists a cluster size weight $p$ such that the FedAvg solution would lead to a worse fairness performance compared to the clustered FL solutions. From Table 10, when the condition $\alpha^{1}_{g}\geq\alpha^{0}_{g}$ holds, there exists a combination of group rates (rows 1-6) such that the FedAvg solution would lead to a worse EqOp fairness performance. These observations from Table 9 and 10 are also consistent with our findings in the Proposition 1 and  2.
E.10 Additional experiments under Gaussian distribution
Simialr to experiments in E.8 and E.9, we now release all settings we imposed before, while we data information for $\mathcal{C}_{\beta}$ unchanged.
Table 11: Cluster $\mathcal{C}_{\alpha}$ SP fairness performance under Gaussian distribution without equalized distance, label rate and group rate
Distribution Label rate Group rate $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{SP}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) $(r_{a},r_{b})$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4.5, 6, 3, 1) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.2598 0.2649 $\uparrow$ 0.2902 $\uparrow$
(0.3, 0.7) 0.2589 0.2593 $\uparrow$ 0.2655 $\uparrow$
(0.7, 0.3) 0.2646 0.2781 $\uparrow$ 0.3074 $\uparrow$
(0.7, 0.3, 0.4, 0.6) (0.5, 0.5) 0.4263 0.4220 $\downarrow$ 0.3917 $\downarrow$
(0.3, 0.7) 0.4288 0.4248 $\downarrow$ 0.3222 $\downarrow$
(0.7, 0.3) 0.4240 0.4198 $\downarrow$ 0.3971 $\downarrow$
(7, 4, 6, 3.5, 1) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.1871 0.2046 $\uparrow$ 0.2483 $\uparrow$
(0.3, 0.7) 0.1785 0.1910 $\uparrow$ 0.2167 $\uparrow$
(0.7, 0.3) 0.1984 0.2236 $\uparrow$ 0.2784 $\uparrow$
(0.7, 0.3, 0.4, 0.6) (0.5, 0.5) 0.3576 0.3752 $\uparrow$ 0.3882 $\uparrow$
(0.3, 0.7) 0.3538 0.3697 $\uparrow$ 0.3335 $\downarrow$
(0.7, 0.3) 0.3620 0.3798 $\uparrow$ 0.3903 $\uparrow$
Table 12: Cluster $\mathcal{C}_{\alpha}$ EqOp fairness performance under Gaussian distribution without equalized distance, label rate and group rate
Distribution Label rate Group rate $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{\alpha})$ $\Delta^{\alpha}_{\texttt{EqOp}}(\theta^{*}_{G})$
($\mu^{1}_{a}$, $\mu^{0}_{a}$, $\mu^{1}_{b}$, $\mu^{0}_{b}$, $\sigma$) ($\alpha^{1}_{a}$, $\alpha^{0}_{a}$, $\alpha^{1}_{b}$, $\alpha^{0}_{b}$) $(r_{a},r_{b})$ $p=\frac{2}{3}$ $p=\frac{1}{2}$
(7, 4.5, 6, 3, 2) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.0993 0.1251 $\uparrow$ 0.1796 $\uparrow$
(0.3, 0.7) 0.0947 0.1115 $\uparrow$ 0.1780 $\uparrow$
(0.7, 0.3) 0.1045 0.1635 $\uparrow$ 0.1814 $\uparrow$
(0.3, 0.7, 0.2, 0.8) (0.5, 0.5) 0.1948 0.1610 $\downarrow$ 0.1558 $\downarrow$
(0.3, 0.7) 0.1967 0.1610 $\downarrow$ 0.1558 $\downarrow$
(0.7, 0.3) 0.1924 0.1615 $\downarrow$ 0.1558 $\downarrow$
(7, 4, 6, 3.5, 2) (0.7, 0.3, 0.6, 0.4) (0.5, 0.5) 0.1051 0.1409 $\uparrow$ 0.1799 $\uparrow$
(0.3, 0.7) 0.1016 0.1293 $\uparrow$ 0.1776 $\uparrow$
(0.7, 0.3) 0.1080 0.1564 $\uparrow$ 0.1822 $\uparrow$
(0.3, 0.7, 0.2, 0.8) (0.5, 0.5) 0.1959 0.1625 $\downarrow$ 0.1569 $\downarrow$
(0.3, 0.7) 0.1950 0.1605 $\downarrow$ 0.1553 $\downarrow$
(0.7, 0.3) 0.1964 0.1650 $\downarrow$ 0.1579 $\downarrow$
From Table 11, we can observe that when the majority of samples are labeled 1 (rows 1-3 and 7-9), there exists a cluster size weight $p$ such that the FedAvg solution would lead to a worse SP fairness performance compared to the clustered FL solutions, which also experimentally extends our findings in the Proposition 2 to the case of an unequalized gap. However, when the majority of samples are labeled differently (rows 4-6 and 10-12), we could find that when $\mu^{1}_{a}-\mu^{0}_{a}>\mu^{1}_{b}-\mu^{0}_{b}$, there exists a $p$ such that the FedAvg solution would lead to a worse SP fairness performance, and a distinct outcome occurs when $\mu^{1}_{a}-\mu^{0}_{a}<\mu^{1}_{b}-\mu^{0}_{b}$. One reason for the distinct behaviors is that the corresponding condition is not satisfied for the experiments in rows 4-6. Additionally, we find that as $p$ enlarges in row 11, the fairness gap decreases, and it could have better fairness performance than using the clustered FL solution. This observation is also consistent with the previous finding that the fairness gap would increase initially and then decrease in the proof of Proposition 2. As we described earlier, it is clearly that for row 11, $p=1/2$ is not in the range of $(p^{\mathcal{C}_{\alpha}}_{low},p^{\mathcal{C}_{\alpha}}_{high})$.
From Table 12, we could observe that when the condition $\alpha^{1}_{g}\geq\alpha^{0}_{g}$ holds (rows 1-3 and 7-9), the changes in the group rates, label rates, and distribution distance do not affect the EqOp fairness performance in the cluster $\mathcal{C}_{\alpha}$. There exists a cluster size weight $p$ such that the FedAvg solution would lead to a worse fairness performance. However, when the condition is not met (rows 4-6 and 10-12), the FedAvg solution would have a better EqOp fairness performance.